Should print the current date \today (in german) on the first two pages, light, outlined. Should print the current date \today (in german) on bottom of the page after the second page. Check that the lines with ˜ ) have correct spacing.

\usepackage[light,firsttwo,outline,bottomafter]{draftcopy}

and

\draftcopyName{\today}{130}
\draftcopySetScale{80}

in the preamble Test10: This is a test page 1 test the tilde tilde tilde (1 fixed space (˜ ) between the words) test the tilde tilde tilde (2 fixed spaces (˜ ) between the words) test the tilde tilde tilde (3 fixed spaces (˜ ) between the words)

This is a testpage 1

Test10: This is a test page 2 test the tilde tilde tilde (1 fixed space (˜ ) between the words) test the tilde tilde tilde (2 fixed spaces (˜ ) between the words) test the tilde tilde tilde (3 fixed spaces (˜ ) between the words)

This is a testpage 2

Test10: This is a test page 3 test the tilde tilde tilde (1 fixed space (˜ ) between the words) test the tilde tilde tilde (2 fixed spaces (˜ ) between the words) test the tilde tilde tilde (3 fixed spaces (˜ ) between the words)

This is a testpage 3

Test10: This is a test page 4 test the tilde tilde tilde (1 fixed space (˜ ) between the words) test the tilde tilde tilde (2 fixed spaces (˜ ) between the words) test the tilde tilde tilde (3 fixed spaces (˜ ) between the words)

This is a testpage 4